OpenStudy (anonymous):

if the equation is an identity,write identity if has no solution rite no solution show work lol 5.5-3b=2b-6.25

OpenStudy (anonymous):

I dont think it's an identity equation but it has a solution which is 2.35

OpenStudy (anonymous):

if it was identity then 5.5=-6.25; absurd.

OpenStudy (radar):

5.5 - 3b = 2b -6.25 subtract 5.5 from both sides -3b= 2b -6.25-5.5 combine and subtract 2b from both sides -5b = -11.75 divide both sides by -5 b=2.35 this is a solution, therefore not an identity.

OpenStudy (radar):

x + y = 8 2x + 2y = 16 2 equations with two unknown, but can't be solved as they are identities.